OpenStudy (dls):

Let f(x) be a non negative continuous function such that the area bounded by the curve y=f(x),x axis and the ordinates x=pi/4 , x=Beta >pi/4 is.... then f(pi/2)=?

OpenStudy (dls):

\[\LARGE Area= (\beta \sin \beta+\frac{\pi}{4} \cos \beta+\sqrt(2)\beta)\]

OpenStudy (dls):

@ganeshie8

OpenStudy (dls):

\[\LARGE \int\limits_\frac{\pi}{4}^\beta f(x)dx=\LARGE (\beta \sin \beta+\frac{\pi}{4} \cos \beta+\sqrt(2)\beta)\] seems so..

ganeshie8 (ganeshie8):

try differentiating \(A(\beta)\), to get \(f(\beta)\) then sub \(\beta = \pi/2\)

OpenStudy (dls):

how? isnt Beta a constant so itll be 0

ganeshie8 (ganeshie8):

\(\beta\) varies from pi/4 to infinity

ganeshie8 (ganeshie8):

int pi/4 to \(\beta\) is same as int 0 to \(x\) if we shift the origin, right ?

ganeshie8 (ganeshie8):